Solve the valid solutions of the following equations providing your values for some of them. Write your answers in solution set form. Identify the extraneous solution, if there is.


logb3 + logb(4x - 5) = 2


(provide a value for b which must be an integer greater than 1)


1
Expert's answer
2021-05-20T17:18:08-0400
"\log_b3+\log_b(4x-5)=2"

"\begin{matrix}
 b>0 \\
 b\not=1\\4x-5>0
\end{matrix}"


"\log_b(3(4x-5))=2"

"3(4x-5)=b^2"

"4x-5=\dfrac{b^2}{3}"

"x=\dfrac{b^2+15}{12}, b>0, b\not=1"



"\big\{\dfrac{b^2+15}{12}\big\},b>0, b\not=1"

There is no the extraneous solution.